OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@mood @apple_pi

OpenStudy (anonymous):

do you know what point slope form is?

OpenStudy (anonymous):

I think I have it y+1=6(x+8)

OpenStudy (anonymous):

you have it backwards switch the 1 and 8

OpenStudy (butterflydreamer):

you're very close! But you mixed up your x and y points that you subbed in :) Remember point slope is: y - y1 = m ( x - x1) So we know slope (m) = 6 , and it passes through the point (-1, -8) So x1 = -1 and y1 = -8 therefore when we sub it into the point-slope formula.. we get: y - ( -8) = 6 ( x - (-1)) y + 8 = 6 ( x + 1)

OpenStudy (anonymous):

Thanks!

OpenStudy (butterflydreamer):

no problem! :)